# Distributed Integrated Sensing, Localization, and Communications over LEO Satellite Constellations

### v1

@@ Abstract
Low Earth orbit (LEO) satellite constellations are rapidly becoming essential enablers of next-generation wireless systems, offering global broadband access, high-precision localization, and reliable sensing beyond terrestrial coverage. However, the inherent limitations of individual LEO satellites, including restricted power, limited antenna aperture, and constrained onboard processing, hinder their ability to meet the growing demands of 6G applications. To address these challenges, this article introduces the concept of distributed integrated sensing, localization, and communication (DISLAC) over LEO constellations, inspired by distributed multiple input multiple output architectures. By enabling inter-satellite cooperation through inter-satellite links, DISLAC can substantially improve throughput, positioning accuracy, and sensing robustness. We present illustrative case studies that quantify these benefits and analyze key system-level considerations, including synchronization, antenna reconfigurability, and ISL design. The article concludes by outlining open research directions to advance the practical deployment of DISLAC in future non-terrestrial networks.
Index Terms:
LEO satellite, NTN, DISAC, DISLAC, 6G.
@@ I Introduction
leo satellite systems are becoming integral to the 6G vision as a key enabler of emerging non-terrestrial networks (NTN), offering broadband connectivity across land, sea, air, and space. Their ability to operate beyond the reach of terrestrial infrastructure helps bridge the digital divide and ensures resilient connectivity during disasters and in remote regions. The technical landscape is evolving accordingly. 3rd generation partnership project (3GPP) has standardized NTN for integration into both enhanced mobile broadband and machine-type communications [1]. Beyond connectivity, low Earth orbit (LEO) satellites now support diverse services including Earth observation, environmental sensing, and independent positioning, navigation, and timing (PNT), complementing or substituting global navigation satellite system (GNSS) [2, 3, 4]. This evolution motivates extending the integrated sensing, localization, and communications (ISLAC) paradigm [5], which integrates sensing, localization, and communication, from terrestrial platforms to space-based architectures, enabling multifunctional LEO systems capable of operating far beyond ground network coverage.
@@
However, individual LEO satellites remain limited by tight power budgets, small antenna apertures, and constrained onboard processing, making it difficult to meet the demands of high-throughput communication and precise sensing or localization. To overcome these limitations, we introduce the concept of distributed ISLAC (DISLAC) over LEO constellations. Drawing from the principles of distributed multiple input multiple output (D-MIMO) in terrestrial systems [6, 5], DISLAC leverages the scale and density of modern constellations and enables cooperation among satellites through inter-satellite links. Such cooperation enables flexible beamforming to boost communication rates and enhances localization and sensing via spatial diversity and fused low- signal-to-noise ratio (SNR) observations. Elements of this vision have been explored in two research threads. One examines co-located ISLACs, where all functionalities are performed on a single satellite using large antenna arrays, but without inter-satellite cooperation [3, 7]. The other investigates distributed LEO constellations addressing sensing, localization, and communication independently [8, 9]. Yet, a unified system-level framework for DISLAC remains missing.
@@
This article aims to fill that gap. As shown in Fig. 1, we consider a cooperative scenario in which multiple multi-antenna LEO satellites jointly perform communication, localization, and sensing with support of ISLs. We analyze the benefits of this architecture through illustrative case studies and explore critical system-level considerations such as synchronization, ISL constraints, antenna flexibility, and scalability. Finally, we outline open research directions to guide the practical development of next-generation DISLAC-enabled NTNs.
@@ II-B Localization Perspective
The distributed nature of LEO satellites enables coordinated multi-satellite localization. In such setups, multiple satellites simultaneously transmit orthogonal waveforms (e.g., separated in frequency, code, or time) while remaining tightly synchronized in time and frequency. The user equipment (UE) can jointly estimate delay, Doppler, or angle-based observables from these signals. Using orthogonal frequency division multiplexing (OFDM) as waveform candidate is natural due to its compatibility with communications and favorable time-frequency resolution, though high Doppler and delay spreads in LEO scenarios must be managed carefully. Techniques such as timing advance and Doppler pre-compensation at the satellite help mitigate these effects.
@@ III Case Studied in LEO constellation-level communication, localization, and sensing
Figure 2: Throughput and signaling overhead comparison: (a) Sum-rate versus number of LEO satellites $S$ under different beamforming schemes; (b) Signaling overhead of collaborative beamforming versus number of UEs $U$, evaluated under Ring and Star topologies with varying satellite number $S$. (E: Edge node, C: Central node).
@@ III-A Case Study 1: Boosting Communication Throughput via Multi-LEO Satellite Collaborative Beamforming
Figures 2(a) and (b) illustrate the sum rate and signaling overhead under different beamforming strategies, evaluated across varying numbers of satellites $S$ and users $U$. We adopt a decentralized weighted minimal mean squared error (WMMSE) algorithm for collaborative beamforming under two typical LEO network topologies: Ring and Star. Unlike centralized approaches, where a central processing unit (CPU) jointly optimizes all beamformers using global CSI, the decentralized schemes perform local optimization at each satellite, relying only on the exchange of intermediate variables with neighboring satellites according to the respective topology [12]. As a baseline, we also consider the single-satellite service (S3) model, where each user is served by only one satellite.
@@ IV-D Standardizations and Regulations
The commercial deployment of DISLAC systems needs also to fulfill standard and regulatory frameworks, starting from spectrum allocation to space debris. Certainly, the carrier frequency and transmission power define the payload specifications within each satellite, but also the coordination between sensing, localization, and communication services. For instance, the adherence to mobile satellite services (MSS) spectrum regulations, defined at International Telecommunication Union (ITU), and co-existence with adjacent or terrestrial services impacts the design of DISLAC payloads. In addition, potential adaptations of the physical layer and signaling optimized for swarm satellites require adoption of novel 3GPP specifications in 6G. Last but not least, space sustainability policies related to space debris and environmental pollution are expected to limit the size of the satellite constellations, leading to further deployment optimizations.
@@ V Emerging Challenges and Research Directions
In addition to the system-level considerations from Table I, several additional challenges must be addressed to realize the full benefits of DISLAC over LEO satellite constellations. A selection of these is provided below.
@@ V-A Collaborative Beamforming over Stochastic LEO Satellite Constellation Topology
Unlike terrestrial networks, where access points are typically fixed with rigid topology, LEO satellite constellations exhibit a dynamic topology that evolves over time due to satellite mobility. Moreover, instead of relying on stable fiber-wired backhaul as in terrestrial systems, LEO satellites are interconnected via ISLs, which are often implemented using highly directional communications (e.g., laser links) that are prone to pointing errors. These unique features render LEO satellite constellations inherently stochastic, making conventional collaborative beamforming, designed for relatively deterministic terrestrial networks, unsuitable for direct application. This necessitates the development of specialized collaborative beamforming schemes that can effectively account for these dynamic impairments and ensure robust performance.
V-B Satellite Selection and Orbit Uncertainties in Localization
An important avenue for future research lies in improving localization performance by selectively incorporating or excluding satellite observations based on their time/phase synchronization and state estimation quality. Not all LEO satellites may maintain accurate knowledge of their own positions and velocities, and/or be accurately synchronized to the satellites network, resulting in poor timing advance and Doppler pre-compensation. In such cases, blindly aggregating data from all available satellites can degrade the localization accuracy of the UE. A more nuanced approach would be to assess the reliability of each satellite in real time and weigh its contribution accordingly. In addition, orbit uncertainties, inherent in LEO constellations due to frequent maneuvers and limited onboard calibration, may pose a significant challenge for localization. However, with dense constellations, these individual errors can be averaged out through statistical averaging. Exploring robust estimators that exploit this averaging effect, while remaining sensitive to outliers introduced by poorly calibrated nodes, is a promising direction for improving the resilience and accuracy of satellite-based localization.
@@ V-C Constellation Optimization and Integration with Terrestrial Systems
The performance of distributed LEO-based sensing and localization strongly depends on constellation geometry and integration with terrestrial infrastructure. For multi-satellite sensing, estimation accuracy, characterized by the Cramér-Rao bound (CRB), is critically influenced by satellite positions. Similarly, wider spatial apertures improve localization accuracy by reducing geometric dilution of precision (GDOP), yet also introduce larger signal delays and Doppler spreads, complicating synchronization and compensation. These trade-offs necessitate careful constellation design and dynamic optimization. Maneuverable modern LEO satellites can adjust constellation geometry based on tasks, balancing wide-area coverage and focused beamforming, particularly for opportunistic sensing such as debris detection, while accounting for energy efficiency, latency, and collision avoidance. Furthermore, integrating LEO-based systems with terrestrial networks, including D-MIMO and reconfigurable intelligent surfaces (RISs), significantly enhances localization in satellite-limited environments. Such integration demands flexible architectures to support robust global localization and sensing through multi-level data fusion.
@@ V-D New Waveforms for Space Sensing with Communication Compatibility
Waveforms critically influence sensing performance through desirable characteristics such as low sidelobes in range-Doppler domains and Doppler resilience. As discussed in Section III-B, robust waveform performance under large Doppler shifts is essential. Although pre- and post-compensation techniques are used in high-Doppler scenarios, their effectiveness remains limited. Frequency-modulated continuous-wave (FMCW) waveforms offer strong Doppler tolerance and simplified receiver, but cannot efficiently support high data rates. Conversely, OFDM waveforms lose subcarrier orthogonality under high Doppler conditions, degrading both communication and sensing performance. This motivates the exploration of communication-compatible waveforms explicitly optimized for sensing. Promising approaches include waveform optimization through problem formulation and emerging waveform schemes such as orthogonal time frequency space (OTFS). While the ambiguity function of OTFS remains under active study, its application to satellite-specific scenarios, particularly in distributed LEO-based multistatic sensing, offers unique opportunities. Carefully optimizing the waveform to balance data rate, sidelobe suppression, and Doppler tolerance will be key to unlocking its full potential in this distributed LEO satellite constellations.
@@ VI Acknowledgments
Yuan Liu (yuan.liu@uni.lu) is a doctoral candidate at the Interdisciplinary Centre for Security, Reliability and Trust (SnT), University of Luxembourg, L-1855, Luxembourg. His current research focuses on channel modeling and signal processing for active sensing applications.
Linlong Wu (linlong.wu@uni.lu) is a Research Scientist at Interdisciplinary Centre for Security, Reliability and Trust (SnT), University of Luxembourg, L-1855, Luxembourg. His current research focuses on signal processing and optimization with applications on radar, ISAC and RIS.
@@
Tareq Y. Al-Naffouri (tareq.alnaffouri@kaust.edu.sa) is a professor in the Electrical and Computer Engineering Program, King Abdullah University of Science and Technology (KAUST), Kingdom of Saudi Arabia. His current research focuses on the areas of sparse, adaptive, and statistical signal processing and their applications, as well as machine learning and network information theory.

### v2

@@ Abstract
Low Earth orbit (LEO) satellite constellations are rapidly becoming essential enablers of next-generation wireless systems, offering global broadband access, high-precision localization, and reliable sensing beyond terrestrial coverage. However, the inherent limitations of individual LEO satellites, including restricted power, limited antenna aperture, and constrained onboard processing, hinder their ability to meet the growing demands of 6G applications. To address these challenges, this article introduces the concept of distributed integrated sensing, localization, and communication (DISLAC) over LEO constellations, inspired by distributed multiple input multiple output architectures. By enabling inter-satellite cooperation through inter-satellite links, DISLAC jointly exploits communication, localization, and sensing functionalities, achieving synergistic gains in throughput, positioning accuracy, and sensing robustness through shared resources and cooperative design. We present illustrative case studies that quantify these benefits and analyze key system-level considerations, including synchronization, antenna reconfigurability, and inter-satellite link design. The article concludes by outlining open research directions to advance the practical deployment of DISLAC in future non-terrestrial networks.
@@ I Introduction
leo satellite systems are becoming integral to the 6G vision as a key enabler of emerging non-terrestrial networks (NTN), offering broadband connectivity across land, sea, air, and space. Their ability to operate beyond the reach of terrestrial infrastructure helps bridge the digital divide and ensures resilient connectivity during disasters and in remote regions. The technical landscape is evolving accordingly. The 3rd generation partnership project (3GPP) has standardized NTN for integration into both enhanced mobile broadband and machine-type communications [1]. Beyond connectivity, low Earth orbit (LEO) satellites now support diverse services including Earth observation, environmental sensing, and independent positioning, navigation, and timing (PNT), complementing or substituting global navigation satellite system (GNSS) [4, 13, 8]. This evolution motivates extending the integrated sensing, localization, and communications (ISLAC) paradigm [9], which integrates sensing, localization, and communication, from terrestrial platforms to space-based architectures, enabling multifunctional LEO systems capable of operating far beyond ground network coverage.
@@
However, individual LEO satellites remain limited by tight power budgets, small antenna apertures, and constrained onboard processing, making it difficult to meet the demands of high-throughput communication and precise sensing or localization. To overcome these limitations, we introduce the concept of distributed ISLAC (DISLAC) over LEO constellations. Drawing from the principles of distributed multiple input multiple output (D-MIMO) in terrestrial systems [3, 9], DISLAC leverages the scale and density of modern constellations and enables cooperation among satellites through inter-satellite links. Such cooperation enables flexible beamforming to boost communication rates and enhances localization and sensing via spatial diversity and fused low- signal-to-noise ratio (SNR) observations. Existing studies on this topic can be broadly grouped into two threads: (i) single-satellite ISLAC systems, where all functionalities are implemented on a single LEO satellite without inter-satellite cooperation [13, 14]; and (ii) distributed LEO constellations, where multiple satellites cooperate but treat communication, localization, and sensing as separate functions [2, 11]. Among these, [11] jointly considers communication and sensing, yet omits user equipment (UE) localization and overlooks key system-level aspects such as synchronization, ISL design, and network topology. Hence, a unified framework that jointly integrates all three functionalities across distributed LEO constellations and captures their mutual interactions remains unexplored.
@@
To the best of our knowledge, this article is the first to present a comprehensive DISLAC framework that unifies communication, localization, and sensing at the constellation level, building on recent advances in distributed integrated sensing and communications (DISAC)/DISLAC for terrestrial D-MIMO networks [3, 9] and extending them to the NTNs. As shown in Fig. 1, we consider a cooperative scenario in which multiple multi-antenna LEO satellites jointly perform communication, localization, and sensing with support of ISLs. We analyze the benefits of this architecture through illustrative case studies and explore critical system-level considerations such as synchronization, ISL constraints, antenna flexibility, and scalability. Finally, we outline open research directions to guide the practical development of next-generation DISLAC-enabled NTNs.
@@ II-B Localization Perspective
The distributed nature of LEO satellites enables coordinated multi-satellite localization. In such setups, multiple satellites simultaneously transmit orthogonal waveforms (e.g., separated in frequency, code, or time) while remaining tightly synchronized in time and frequency. The UE can jointly estimate delay, Doppler, or angle-based observables from these signals. Using orthogonal frequency division multiplexing (OFDM) as waveform candidate is natural due to its compatibility with communications and favorable time-frequency resolution, though high Doppler and delay spreads in LEO scenarios must be managed carefully. Techniques such as timing advance and Doppler pre-compensation at the satellite help mitigate these effects.
@@ II-C Sensing Perspective
II-D Mutual Benefits and Trade-offs
In a unified DISLAC framework, communication, localization, and sensing become mutually reinforcing rather than isolated functionalities, as conceptually illustrated in Fig. 2. Accurate UE positioning and network-wide LEO localization can significantly improve geometry-dependent channel estimation, enable position-aware beamforming, and support efficient user association, scheduling, and handover management. Conversely, high-throughput service links and ISLs provide the data backbone for distributing sensing observations and localization coordination across the constellation, as well as for fusing multi-LEO sensing data. Sensing measurements, in turn, enhance situational awareness by detecting obstacles, atmospheric disturbances, or debris, thereby supporting adaptive link management and robust service continuity. These synergies also introduce inherent resource trade-offs: spectrum and waveform resources must be shared between throughput and positioning accuracy, sensing resolution competes with communication throughput, and spatial beam patterns need to be allocated between illuminating UEs for localization and illuminating passive targets for sensing. Balancing these coupled benefits and trade-offs calls for flexible resource orchestration and adaptive ISL coordination, highlighting the distinctive system-level design challenges of integrating communication, localization, and sensing in distributed LEO constellations.
Figure 2: Conceptual illustration of exemplary synergies and trade-offs among communication, localization, and sensing in a unified DISLAC framework over LEO constellations. “Communication” aggregates service links between LEOs and UEs and ISLs; “Localization” covers UE positioning and network-wide satellite localization; “Sensing” includes ground and inter-LEO sensing. Blue arrows illustrate example mutual benefits, such as position-aware beamforming and handover, high-throughput links for localization coordination, ISL-enabled fusion of sensing data, and shared spectrum/pilots/synchronization for joint situational awareness. Orange arcs indicate representative resource trade-offs in spectrum/throughput versus positioning accuracy, throughput demand versus sensing resolution, and spatial beam allocation between UE localization and passive target sensing.
@@ III Case Studied in LEO constellation-level communication, localization, and sensing
Figure 3: Throughput and signaling overhead comparison: (a) Sum-rate versus number of LEO satellites $L$ under different beamforming schemes; (b) Signaling overhead of collaborative beamforming versus number of UEs $U$, evaluated under Ring and Star topologies with varying satellite number $L$. (E: Edge node, C: Central node).
@@ III-A Case Study 1: Boosting Communication Throughput via Multi-LEO Satellite Collaborative Beamforming
Figures 3(a) and (b) illustrate the sum rate and signaling overhead under different beamforming strategies, evaluated across varying numbers of satellites $L$ and users $U$. We adopt a decentralized weighted minimal mean squared error (WMMSE) algorithm for collaborative beamforming under two typical LEO network topologies: Ring and Star. Unlike centralized approaches, where a central processing unit (CPU) jointly optimizes all beamformers using global CSI, the decentralized schemes perform local optimization at each satellite, relying only on the exchange of intermediate variables with neighboring satellites according to the respective topology [15]. As a baseline, we also consider the single-satellite service ($\text{S}^{3}$) model, where each user is served by only one satellite.
@@ IV DISLAC Over LEO Satellite Constellation — A System-Level Perspective
The case studies in Section III have illustrated how multi-LEO cooperation influences each functionality individually. In a unified DISLAC framework, these functionalities become mutually reinforcing. On the one hand, accurate sensing and localization strengthen communication through more reliable user association, seamless handover, and efficient resource management. On the other hand, advanced communication capabilities further enhance sensing and localization by enabling high-quality signal transmission, flexible information exchange, and greater spatial diversity. Section IV therefore builds on the observations from Section III and focuses on the NTN-specific, system-level aspects of DISLAC integration, such as synchronization, ISLs, constellation topology, and regulatory constraints, rather than repeating numerical comparisons for a single joint experiment.
@@ IV-D Standardizations and Regulations
The practical deployment of DISLAC systems must comply with existing and emerging standardization and regulatory frameworks, ranging from spectrum management to orbital sustainability. In particular, the legacy approach asking each communication, navigation or sensing system operator to comply with the regulations established at International Telecommunication Union (ITU) for a mobile satellite service (MSS), radio navigation-satellite service (RNSS) or Earth exploration satellite service (EESS), needs to be addressed when integrating all three services in a single system. At first glance, a joint compliance over all services could further constrain the DISLAC payload and network design. However, DISLAC also represents an opportunity to alleviate such burden, building on the exploitation of multi-purpose signals, which revisits the notion of primary or secondary services (i.e., co-existing in the same band) depending on the use of those signals. Furthermore, realizing inter-satellite cooperation at scale will require extensions to current 3GPP NTN specifications to support standardized inter-satellite signaling, synchronization, and multi-function coexistence. Beyond spectrum policy, space sustainability and debris-mitigation regulations impose additional constraints on constellation size, orbit selection, and resource reuse, urging more efficient and environmentally responsible deployment strategies. By integrating these technical and regulatory considerations, DISLAC provides a forward-looking framework that links system-level innovation with the industrial standardization pathways essential for large-scale multi-satellite collaboration.
@@ V Emerging Challenges and Research Directions
In addition to the system-level considerations summarized in Table I, several additional challenges must be addressed to realize the full benefits of DISLAC over LEO satellite constellations. The following subsections discuss key research directions from a DISLAC perspective.
@@ V-A Collaborative Beamforming over Stochastic LEO Satellite Constellation Topology
Unlike terrestrial networks, where access points are typically fixed with rigid topology, LEO satellite constellations exhibit a dynamic topology that evolves over time due to orbital motion. Moreover, instead of relying on stable fiber-wired backhaul as in terrestrial systems, LEO satellites are interconnected via ISLs, often implemented using highly directional links (e.g., optical), which are prone to pointing errors and link interruptions. These features render LEO constellations inherently stochastic, making conventional collaborative beamforming schemes, developed for deterministic terrestrial networks, unsuitable for direct application. Future work should therefore explore topology-aware and synchronization-aided beamforming strategies that jointly exploit localization and sensing information to improve link alignment, mitigate Doppler uncertainty, and enhance multi-LEO coordination robustness within the DISLAC framework.
V-B Satellite Selection and Orbit Uncertainties
An important research direction lies in improving localization performance by selectively incorporating or excluding satellite observations based on their time/phase synchronization and state-estimation quality. Not all LEO satellites may maintain accurate knowledge of their own positions and velocities or remain tightly synchronized within the constellation, resulting in imperfect timing advance and Doppler pre-compensation. In such cases, blindly aggregating data from all available satellites can degrade localization accuracy. A more nuanced approach is to assess each satellite’s reliability in real time and weigh its contribution accordingly. Orbit and state uncertainties, inherent to LEO constellations due to frequent maneuvers and limited onboard calibration, can also propagate to sensing and communication functions through synchronization and beam alignment errors. Hence, developing joint reliability metrics and cross-functional calibration methods will be essential to ensure robust and mutually consistent localization, sensing, and communication performance across the constellation.
@@ V-C Constellation Optimization and Integration with Terrestrial Systems
The performance of distributed LEO-based sensing and localization strongly depends on constellation geometry and its integration with terrestrial infrastructure. For multi-satellite sensing, estimation accuracy, characterized by the Cramér-Rao bound (CRB), is critically influenced by satellite positions. Similarly, wider spatial apertures improve localization accuracy by reducing geometric dilution of precision (GDOP), yet also introduce larger signal delays and Doppler spreads, complicating synchronization and waveform reuse across communication and sensing functions. These trade-offs necessitate constellation designs that adapt geometry and resource allocation according to joint service requirements. Maneuverable modern LEO satellites can dynamically adjust their geometry to balance wide-area coverage and focused beamforming, enabling task-driven reconfiguration for both data and sensing missions. Furthermore, integrating LEO-based systems with terrestrial D-MIMO and reconfigurable intelligent surfaces (RISs) offers additional cross-domain diversity and resource sharing. Such integration demands flexible architectures capable of fusing multi-source information for global-scale DISLAC operation.
@@ V-D New Waveforms for Space Sensing with Communication Compatibility
Waveforms critically influence sensing and communication performance through characteristics such as low sidelobes, Doppler resilience, and spectral efficiency. As discussed in Section III-B, robust waveform behavior under large Doppler shifts is essential in LEO environments. Although pre- and post-compensation techniques are commonly employed, their effectiveness remains limited under fast-varying dynamics. Frequency-modulated continuous-wave (FMCW) waveforms offer strong Doppler tolerance and simplified receivers but cannot efficiently support high data rates. Conversely, OFDM waveforms provide high throughput yet lose subcarrier orthogonality under large Doppler spreads, degrading both communication and sensing performance. These contrasting properties motivate the design of communication-compatible sensing waveforms that balance data rate, ambiguity resolution, and Doppler tolerance, such as orthogonal time frequency space (OTFS)-based designs. Developing adaptive waveform strategies that jointly optimize communication throughput, localization precision, and sensing accuracy will be key to fully realizing the potential of DISLAC systems.
@@ VI Acknowledgments
Yuan Liu (yuan.liu@uni.lu) is a Postdoctoral Researcher at the Interdisciplinary Centre for Security, Reliability and Trust (SnT), University of Luxembourg, L-1855, Luxembourg. His current research focuses on channel modeling and signal processing for active sensing applications.
Linlong Wu (linlong.wu@uestc.edu.cn) is a professor with the School of Information and Communication Engineering, University of Electronic Science and Technology of China (UESTC), Chengdu, China. His research focuses on signal processing, radar, ISAC and RIS.
@@
Tareq Y. Al-Naffouri (tareq.alnaffouri@kaust.edu.sa) is a professor in the Electrical and Computer Engineering Program, King Abdullah University of Science and Technology (KAUST), Kingdom of Saudi Arabia. His current research focuses on statistical inference and learning, with applications in wireless communications, positioning, navigation, smart cities, and smart health.